OpenStudy (anonymous):

Help Please! <3 Tell whether or not the given point is a solution to the system of equations. (5,3) y=-0.75x+0.75 y=-2/3x+1/3

OpenStudy (ivettef365):

Just try the points to see if is true, you have (5,3) x=5 and y = 3 therefore on the first equation 3 = 0.75(5) + 0.75 3 = 3.75 + 0.75 3 = 4.5 this is not true therefore is not a solution

OpenStudy (ivettef365):

Sorry, didnt see the (-) sign

OpenStudy (ivettef365):

then you have 3 = -0.75(5) + 0.75 3 = - 3.75 + 0.75 3 = -3 (still not true)

OpenStudy (anonymous):

Thank you so much!! GBU!! :)